Cody’s car accelerates from 0m/s to 45 m/s northward in 15 seconds. What is

the acceleration of the car?


1
Expert's answer
2022-04-14T11:18:36-0400

The acceleration of the car

"a=\frac{\Delta v}{\Delta t}=\rm\frac{45\: m/s-0\: m/s}{15\: s}=3\: m/s^2"

The acceleration is directed northward.